drayasdfghjkl:

What is 125x9 + 64y12 written as a sum of cubes?

TheSmartOne:

Here's the formula a^3 + b^3 = (a+b)(a^2-ab+b^2) So what are your a and b values? What cubed will get you 125 and 64? You don't have to try very hard, just start from 2^3 and go up a few numbers Just a recap if you forgot how exponents work 2^3 = 2*2*2 = 8